A 12-kg box is released from the top of an incline that is 5.0m long and makes an angle of 40° to the horizontal. A 60-N friction force impedes the motion of the box. (a) What will be the acceleration of the box? (b) How long will it take to reach the bottom of the incline?




1
Expert's answer
2021-12-02T10:03:31-0500

(a)


"mg\sin40°-60=ma\to a=g\sin40°-60/m=9.8\cdot\sin40°-60/12=1.3\ (m/s^2)"


(b)


"l=at^2/2\to t=\sqrt{2l/a}=\sqrt{2\cdot5/1.3}=2.77\ (s)"